OpenStudy (anonymous):

Write this relationship as a product of powers One million is 1000 times as great as one thousand

OpenStudy (mathstudent55):

10^1 = 10 = 10 10^2 = 10 x 10 = 100 10^3 = 10 x 10 x 10 = 1,000 10^4 = 10 x 10 x 10 x 10 = 10,000 10^5 = 10 x 10 x 10 x 10 x 10 = 100,000 10^6 = 10 x 10 x 10 x 10 x 10 x 10 = 1,000,000 Does this help?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

So would it be like 10^6=10^3x10^3?

OpenStudy (mathstudent55):

you got it

OpenStudy (anonymous):

Ok cool, but now I have to do a quotient of powers and it doesn't make sense! One hundred is one-tenth of one thousand

OpenStudy (mathstudent55):

100 = (1/10)(1000) You need to do it with powers also?

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

Remember what I wrote above with 10, 100 etc and the powers of 10?

OpenStudy (anonymous):

Yup

OpenStudy (mathstudent55):

You can keep on going to 10^7, 10^8, etc., but you can also go the other way.

OpenStudy (mathstudent55):

10^2 = 10 x 10 = 100 10 = 10 = 10 10^-1 = 1/10 = 1/10 10^-2 = 1/10^2 = 1/100 10^-3 = 1/10^3 = 1/1,000 etc.

OpenStudy (mathstudent55):

You need to write 100 = (1/10)1,000 100 = 10^2 1/10 = 1/10^-1 1,000 = 10^3

OpenStudy (anonymous):

ok... I kind of understand What about one is one-millionth of one million

OpenStudy (mathstudent55):

one millionth is 1/1,000,000 = 1/10^6 one million is 1,000,000 = 10^6

OpenStudy (anonymous):

Oooh ok

OpenStudy (mathstudent55):

Do you know how to multiply powers with the same base?

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

For example, what is 7^3 x 7^5? without figuring out what each power is

OpenStudy (anonymous):

add 3 and 5 so it's 7^8

OpenStudy (mathstudent55):

Ok, since you know that, now look at that example, what is a millionth of a million? Without thinkiong of thise problem and powers, just answer, what is a millionth of a million?

OpenStudy (anonymous):

0?

OpenStudy (mathstudent55):

No. If you have a million and take one millionth of it? Think of a similar example with a smaller number. What is a tenth of ten?

OpenStudy (anonymous):

1

OpenStudy (mathstudent55):

right. and a millionth of a million is also one.

OpenStudy (anonymous):

Oh!

OpenStudy (mathstudent55):

Now think of the example, one millionth of a million using powers of ten (1/1,000,000)(1,000,000), but we know that 1/1,000,000 is 10^-6 and 1,000,000 is 10^6, so (1/1,000,000)(1,000,000) = 10^-6 x 10^6 Using multiplication of powers, 10^-6 x 10^6 = 10^0 = 1 So you see that one millionth of a million is 1whether you do it with numbers or powers of ten

OpenStudy (anonymous):

Omg thanks!